# ZeroPur: Succinct Training-Free Adversarial Purification

### v1

@@ Abstract
Adversarial purification is a kind of defense technique that can defend various unseen adversarial attacks without modifying the victim classifier. Existing methods often depend on external generative models or cooperation between auxiliary functions and victim classifiers. However, retraining generative models, auxiliary functions, or victim classifiers relies on the domain of the fine-tuned dataset and is computation-consuming. In this work, we suppose that adversarial images are outliers of the natural image manifold and the purification process can be considered as returning them to this manifold. Following this assumption, we present a simple adversarial purification method without further training to purify adversarial images, called ZeroPur. ZeroPur contains two steps: given an adversarial example, Guided Shift obtains the shifted embedding of the adversarial example by the guidance of its blurred counterparts; after that, Adaptive Projection constructs a directional vector by this shifted embedding to provide momentum, projecting adversarial images onto the manifold adaptively. ZeroPur is independent of external models and requires no retraining of victim classifiers or auxiliary functions, relying solely on victim classifiers themselves to achieve purification. Extensive experiments on three datasets (CIFAR-10, CIFAR-100, and ImageNet-1K) using various classifier architectures (ResNet, WideResNet) demonstrate that our method achieves state-of-the-art robust performance. The code will be publicly available.
@@ 1 Introduction
Recent studies show that adding carefully crafted, imperceptible perturbations to natural examples can easily fool deep neural networks (DNNs) to make wrong decisions [42, 11]. This potential vulnerability underlying their remarkable performance raises a significant challenge for security-critical applications. Thus, designing efficient adversarial defense techniques is necessary for real-world applications in DNNs.
@@
Different from adversarial training, adversarial purification [51, 38, 30, 29] aims to remove adversarial perturbations in adversarial examples. These methods do not require adversarial examples in the model training and effectively defend against unseen attacks, making them more applicable in real-world scenarios. However, the existing purification methods often depend on external generative models [10, 40] or cooperation between external auxiliary functions and the victim classifier [38, 29, 19]. Retraining generative models, parameterized auxiliary functions or the classifier relying on the domain of the fine-tuned dataset is computationally demanding and restricts their flexibility.
Inspired by the natural image manifold hypothesis, we suppose adversarial images are outliers of the natural image manifold and the purification process is to return them to this manifold. We present a simple adversarial purification method named ZeroPur, where "Zero" means our method does not need to train any external models, parameterized auxiliary functions, and victim classifiers. As illustrated in Fig. 1, given the adversarial image, ZeroPur comprises two stages, i.e., Guided Shift (GS) and Adaptive Projection (AP), to purify the given image. In detail, GS guides adversarial examples towards the blurred counterparts to target natural images on the manifold to obtain the shifted embeddings. This is achieved by iteratively pulling the distance between adversarial examples and a set of their blurred counterparts in the embedding space. Since the inherent limitations of the low-quality embeddings from blurred counterparts, adversarial examples may not precisely return to the target locations, but these blurred counterparts do provide a reasonable direction. After that, we introduce AP to construct the directional vectors based on these shifted embeddings which provide a reference direction of the movement of adversarial images. The adversarial images are projected adaptively onto the manifold by projection maximization in this direction. Despite its simplicity, ZeroPur consistently outperforms most state-of-the-art adversarial training and purification methods.
@@
  • We present a succinct adversarial purification approach named ZeroPur including two stages: Guided Shift and Adaptive Projection, which requires no retraining of external models, parameterized functions, and the victim classifier.
@@ 2 Review of Literature
External model-based adversarial purification (EBP). Samangouei et al. [35] propose defense-GAN, a generator that models the distribution of natural images, which enables the transformation from adversarial examples to natural images. Song et al. [39] assume that adversarial examples primarily reside in the low probability density region of the training distribution, and design PixelDefend to approximate this distribution using the PixelCNN [44]. Recently, utilizing score-based models [51] and diffusion models [45, 30] as purification models is introduced and shown to achieve significantly improved robust performance. More recently, Lin et al. [26] propose a framework called AToP that fine-tunes the purification model adversarially to integrate the benefits of both adversarial training and adversarial purification. However, these works rely on external generative models, substantial datasets (e.g., crafting adversarial examples) and computational resources.
Auxiliary-based adversarial purification (ABP). In contrast to the above two approaches, ABP tends to introduce an auxiliary function to cooperate with the classifier to purify adversarial images. Shi et al. [38] propose a lightweight purification method SOAP, which uses self-supervised loss to realize online purification. SOAP no longer depends on generative models but requires classifiers to incorporate the corresponding auxiliary loss (self-supervised loss) in the training stage. To further reduce the time cost of purification, recent works introduce parameterized auxiliary functions, allowing purification to be accomplished through the exclusive training of these functions. Such auxiliary functions are designed to be lightweight compared to classifiers so that they can significantly reduce the time cost. For example, Mao et al. [29] introduce a two-layer network that estimates the contrastive features, which can purify the adversarial images. Hwang et al. [19] propose AID-Purifier, an auxiliary discriminator based on information maximization principles that can transform adversarial images into natural images.
@@ 3.1 Adversarial Purification in The Natural Image Manifold
Following the natural image manifold hypothesis [17], natural images are assumed to reside on a specific manifold known as the natural image manifold. Given the classification loss function $\ell$, the learning process of DNNs finding optimal parameters $\theta$ can be formulated by:
@@
where $\epsilon$ is the maximal norm which defines the set of allowed perturbations for a given example $\bm{x}$. In most contexts, such as The Projected Gradient Descent (PGD) [28], $\bm{\delta}^{*}$ is approximated by the local worst-cast $\bm{\delta}$.
@@
For notational simplicity, we denote $f_{1...l}(\bm{x})=f(\bm{x})$ and $e_{l}(\bm{\delta})=e(\bm{\delta})$ when $l$ is the last layer of $f$. Theorem 1 describes how the perturbation $\bm{\delta}$ is amplified layer by layer during the forward of $f$ so that the embedding of natural image $\bm{x}$ deviates manifold $\mathcal{M}$ and become the outlier, thereby demonstrates the reason for the incorrect decision made by $g$:
@@
where $e(\bm{\delta})$ can be viewed as the deviation distance from the natural image manifold.
By this formulation, adversarial purification can be naturally seen as the process of estimating $e({\bm{\delta}})$ in the embedding space, the approximation of the natural image manifold. This means the adversarial embedding $f(\bm{x}_{\mathrm{adv}})=f(\bm{x}+\bm{\delta})$ can return the manifold along the same path $-e(\bm{\delta})$. While precise estimation of $e(\bm{\delta})$ is challenging, it is fortunate that only an approximation, denoted as $\tilde{e}(\bm{\delta})$ is sufficient to move the adversarial embedding back within the manifold.
Figure 2: (a) An illustration of activation of crucial semantic features on CIFAR-10 and ImageNet-1K. (b) Accuracy of different properties of images on three models, where the blurring operator is the median filter. Left: AutoAttack results on CIFAR-10. Middle and Right: AutoAttack on ImageNet.
@@ 3.2 Guided Shift
where $\bm{\delta}_{\mathrm{pfy}}$ and $\epsilon_{\mathrm{pfy}}$ are defined to correspond to $\bm{\delta}$ and $\epsilon$ in Eq.(2) to offset the perturbation. However, even if $\epsilon_{\mathrm{adv}}$ can be considered as a hyperparameter, the ground-truth label $y$ is not accessible. External model-based adversarial purification [35, 51, 17, 30] typically train a purification model to minimize the global $\ell_{\mathrm{pfy}}$. Auxiliary function-based adversarial method [29, 38, 19] tend to design a suitable $\ell_{\mathrm{pfy}}$ to complete purification without relying on external generative models. They need to retrain the classifier or parameterized auxiliary function to effectively cooperate in removing adversarial perturbations, which means these methods all introduce parameters $\Theta$ in $\ell_{\mathrm{pfy}}$:
@@
Based on our assumption in subsection 3.1 that adversarial images are outliers of the natural image manifold and the purification process can be considered as returning them to this manifold, we aim to design $\ell_{\mathrm{pfy}}$ without $\Theta$ by allowing them to shift towards the manifold adaptively:
@@
We are starting to investigate whether a simple image transformation, such as color jitter, grayscale, Gaussian blur, solarization, and equalization can shift adversarial images towards the natural image manifold to improve classifier performance. We find that blurring adversarial examples can improve the robust accuracy of victim classifiers. Additionally, as shown in Fig 2(a), the blur operation can reactivate regions disrupted by adversarial images. This phenomenon suggests that a simple blurring operator can bring adversarial examples closer to the manifold and reverse adversarial properties.
@@
where $\mathrm{blur}(\cdot)$ is the blurring operator and $\bm{x}_{\mathrm{g}}^{0}:=\bm{x}_{\mathrm{adv}}$, such as the median filter, Gaussian blur, and so on. The workflow of GS is shown in Algorithm 1, where we use $\mathrm{sgn}(\cdot)$ to regulate the step size.
To demonstrate the process of GS, we employ two blurring operators: the median filter and Gaussian blur to compute the cosine similarity between single-step result $\bm{x}_{\mathrm{g}}^{t}$ of GS and the natural image $\bm{x}$ ($\star$-$GS$-$N$), as well as between the blurred counterpart $\bm{\tilde{x}}_{\mathrm{g}}^{t}$ of $\bm{x}_{\mathrm{g}}^{t}$ and $\bm{x}$ ($\star$-$GS^{\prime}$-$N$). As shown in Fig. 3, in each step, the cosine similarity ($\star$-$GS$-$N$) is consistently smaller than the cosine similarity $\star$-$GS^{\prime}$-$N$. Furthermore, these cosine similarities consistently increase as GS progresses, and adversarial examples are gradually getting closer to the natural image manifold guided by their blurred counterparts.
@@ 3.3 Adaptive Projection
We also can observe an interesting phenomenon in Fig. 3. Regardless of the type of blurring operators, dash lines ($\star$-$GS^{\prime}$-$N$) and solid lines ($\star$-$GS$-$N$) tend to converge. Therefore, we can have the following assumption:
@@
(a) AP w/o. RP in Adv $\bm{x}_{\mathrm{init}}$
(b) AP in Adv $\bm{x}_{\mathrm{init}}$
(c) AP w/o. RP in Nat $\bm{x}_{\mathrm{init}}$
(d) AP in Nat $\bm{x}_{\mathrm{init}}$
Figure 4: Intutive explanation of AP in the context of adversarial or natural $\bm{x}_{\mathrm{init}}$. PR: the perceptual regularization (Eq.(20)).
@@
$\mathcal{L}_{l}(\bm{x}_{\mathrm{init}},\bm{x}_{\mathrm{g}},\bm{x}_{\mathrm{p}})=-\Delta\bm{u}^{l}_{\mathrm{p}}\cdot\Delta\bm{u}^{l}_{\mathrm{g}},$ (16)
@@
where $f_{l}$ denoted as feature maps at layer $l$ of the classifier, $\bm{x}_{\mathrm{p}}$ is initialized by $\bm{x}_{\mathrm{init}}$. Maximizing Eq.(16) is equivalent to maximizing the projection of $\bm{u}^{l}_{\mathrm{p}}$ onto $\bm{u}^{l}_{\mathrm{g}}$ since $\|\bm{u}^{l}_{\mathrm{g}}\|$ is a constant. The increase in projection implies that $\bm{x}_{\mathrm{p}}$ is not restricted by the blurred images to continue moving along the direction of the result of GS, which allows $\bm{x}_{\mathrm{p}}$ move independently toward the natural image manifold.
However, We typically do not know whether an image is adversarial in real-world applications (i.e., $\bm{x}_{\mathrm{init}}=\bm{x}$ when $\bm{x}$ is natural), repeatedly guiding images by blurred counterparts may cause natural images to deviate from the manifold. As shown in Fig. 4, this deviation will be enhanced by AP. Therefore, we introduce a perceptual regularization term in Eq.(15). Let $F_{1}$ is the dynamic of Eq.(16):
$F_{1}=\nabla_{\bm{x}_{\mathrm{p}}}\frac{1}{\|S\|}\sum_{l\in S}\mathcal{L}_{l}(\bm{x}_{\mathrm{init}},\bm{x}_{\mathrm{g}},\bm{x}_{\mathrm{p}})\ ,$ (18)
@@ 4 Experiments
Datasets and Metrics. Three benchmarks CIFAR-10 [22], CIFAR-100, and ImageNet-1K [9] are considered to evaluate our method. We compare our method with the state-of-the-art adversarial training methods reported in standard benchmark RobustBench [7] and other adversarial purification methods [38, 29, 19, 41, 16, 51, 43, 30, 17, 26, 45]. In all experiments, we consider two metrics to evaluate the performance of all methods: standard accuracy and robust accuracy. The standard accuracy measures the performance of the defense method on natural images, while the robust accuracy measures the performance on adversarial images.
Adversarial attacks and victim classifiers. We evaluate our methods with three attacks: AutoAttack [6], DI2-FGSM [49], and BPDA [3]. AutoAttack is a powerful adaptive attack commonly used in most defense studies. We also use DI2-FGSM to seek whether an attack robust to the blurring operator can affect our method. Additionally, we consider BPDA to attack our purification module in the worst-case scenario. These attacks will attack three victim classifiers including ResNet-18 [14], ResNet-50, and WideResNet-28-10 [53].
Table 1: Standard and robust accuracy (%) against AutoAttack $\ell_{\infty}(\epsilon=8/255)$ in comparison with AT &\&& ABP methods. The first section corresponds to AT methods and the second to ABP methods. $\dagger$ Since Hwang et al. [19] only reports the performance of WideResNet-34-10 in the original paper, and the pre-trained parameterized auxiliary checkpoint has been removed, we use this result for comparison.
Require Training ResNet-18 WideResNet-28-10
Classifier Auxiliary Method Standard Robust Method Standard Robust
@@
(Hwang et al., 2023$\dagger$ 87.02 56.63 (Hwang et al., 2023) 87.02 56.63
@@
(Hwang et al., 2023) 64.73 32.86 (Hwang et al., 2023) 64.73 32.86

### v2

@@ Abstract
Adversarial purification is a kind of defense technique that can defend against various unseen adversarial attacks without modifying the victim classifier. Existing methods often depend on external generative models or cooperation between auxiliary functions and victim classifiers. However, retraining generative models, auxiliary functions, or victim classifiers relies on the domain of the fine-tuned dataset and is computation-consuming. In this work, we suppose that adversarial images are outliers of the natural image manifold, and the purification process can be considered as returning them to this manifold. Following this assumption, we present a simple adversarial purification method without further training to purify adversarial images, called ZeroPur. ZeroPur contains two steps: given an adversarial example, Guided Shift obtains the shifted embedding of the adversarial example by the guidance of its blurred counterparts; after that, Adaptive Projection constructs a directional vector by this shifted embedding to provide momentum, projecting adversarial images onto the manifold adaptively. ZeroPur is independent of external models and requires no retraining of victim classifiers or auxiliary functions, relying solely on victim classifiers themselves to achieve purification. Extensive experiments on three datasets (CIFAR-10, CIFAR-100, and ImageNet-1K) using various classifier architectures (ResNet, WideResNet) demonstrate that our method achieves state-of-the-art robust performance. The code will be publicly available.
@@ 1 Introduction
Over the past decade, a substantial body of work has confirmed that deep neural networks (DNNs) are highly vulnerable to imperceptible adversarial perturbations that cause wrong decisions [47, 10, 32, 4, 54]. This potential vulnerability, despite their remarkable performance, poses a significant challenge for security-critical applications. Thus, designing effective and efficient adversarial defense techniques is essential to ensure the reliability and trustworthiness of DNNs in real-world scenarios.
@@
Different from adversarial training, adversarial purification [56, 42, 33, 31] aims to remove adversarial perturbations in adversarial examples. These methods do not require adversarial examples in the model training and effectively defend against unseen attacks, making them more applicable in real-world scenarios. However, the existing purification methods often depend on external generative models [9, 45] or cooperation between external auxiliary functions and the victim classifier [42, 31, 17]. Retraining generative models, parameterized auxiliary functions, or the classifier relying on the domain of the fine-tuned dataset is computationally demanding and restricts their flexibility.
Inspired by the natural image manifold hypothesis, we suppose that adversarial images are outliers of the natural image manifold, and the purification process aims to return them to this manifold. We propose a simple adversarial purification method called ZeroPur, where ’Zero’ means our method does not need to train any external models, parameterized auxiliary functions, and victim classifiers. As illustrated in Figure 1, ZeroPur comprises two stages: Guided Shift (GS) and Adaptive Projection (AP) for purifying the adversarial image. Specifically, we observe that the blurring operator provides a reasonable direction towards natural image on the manifold for an adversarial sample. Therefore, GS iteratively guides adversarial sample towards its blurred counterparts, obtaining shifted embedding that approximate the manifold. Due to the inherent limitations of low-quality embeddings from blurred counterparts, adversarial sample may not precisely return to the target location, but the shifted embedding after multiple iterations provide a more reliable direction. Subsequently, we introduce AP to construct a directional vector based on this shifted embedding and project the adversarial image adaptively onto the manifold by maximizing the projection along this direction. Despite its embarrassingly simple design, ZeroPur consistently outperforms most state-of-the-art (SOTA) adversarial training and purification methods.
@@
  • We present a succinct adversarial purification approach named ZeroPur, including two stages: Guided Shift and Adaptive Projection, which requires no retraining of external models, parameterized functions, and the victim classifier.
@@ 2 Related Work
External model-based adversarial purification (EBP). Samangouei et al. [39] propose defense-GAN, a generator that models the distribution of natural images, which enables the transformation from adversarial examples to natural images. Song et al. [44] assume that adversarial examples primarily reside in the low probability density region of the training distribution, and design PixelDefend to approximate this distribution using the PixelCNN [49]. Later, utilizing score-based models [56] and diffusion models [43, 33, 25, 18] as purification models is introduced and shown to achieve significantly improved robust performance. Recently, Lin et al. [27] propose a framework called AToP that fine-tunes the purification model adversarially to integrate the benefits of both adversarial training and adversarial purification. However, these works rely on external generative models, substantial datasets (e.g., crafting adversarial examples), and computational resources.
Auxiliary-based adversarial purification (ABP). In contrast to the above two approaches, ABP tends to introduce an auxiliary function to cooperate with the classifier to purify adversarial images. Shi et al. [42] propose a lightweight purification method, SOAP, which uses self-supervised loss to realize online purification. SOAP no longer depends on generative models but requires classifiers to incorporate the corresponding auxiliary loss (self-supervised loss) in the training stage. To further reduce the time cost of purification, recent works introduce parameterized auxiliary functions, allowing purification to be accomplished through the exclusive training of these functions. Such auxiliary functions are designed to be lightweight compared to classifiers so that they can significantly reduce the time cost. For example, Mao et al. [31] introduces a two-layer network that estimates the contrastive features, which can purify the adversarial images. Hwang et al. [17] propose AID-Purifier, an auxiliary discriminator based on information maximization principles that can transform adversarial images into natural images.
@@ 3.1 Adversarial Purification in the Natural Image Manifold
Following the natural image manifold hypothesis [29, 44], natural images are assumed to reside on a specific manifold known as the natural image manifold. Given the classification loss function $\ell$, the learning process of DNNs finding optimal parameters $\theta$ can be formulated by:
@@
where $\epsilon$ is the maximal norm which defines the set of allowed perturbations for a given example $\bm{x}$, in most contexts, such as the Projected Gradient Descent (PGD) [30], $\bm{\delta}^{*}$ is approximated by the local worst-cast $\bm{\delta}$.
@@
The proof of Theorem 1 can be found in Appendix A.1. For notational simplicity, we denote $f_{1...l}(\bm{x})=f(\bm{x})$ and $e_{l}(\bm{\delta})=e(\bm{\delta})$ when $l$ is the last layer of $f$. Theorem 1 describes how the perturbation $\bm{\delta}$ is amplified layer by layer during the forward of $f$ so that the embedding of natural image $\bm{x}$ deviates manifold $\mathcal{M}$ and become the outlier, thereby demonstrates the reason for the incorrect decision made by $g$:
@@
where $e(\bm{\delta})$ can be viewed as the deviation distance from the natural image manifold. As shown in Figure 2(a), the adversarial images induced by malicious attacks on natural samples seriously deviate from the natural image manifold.
By this formulation, adversarial purification can be naturally seen as the process of estimating $e({\bm{\delta}})$ in the embedding space, the approximation of the natural image manifold. This means the adversarial embedding $f(\bm{x}_{\mathrm{adv}})=f(\bm{x}+\bm{\delta})$ can return the manifold along the same path $-e(\bm{\delta})$. While precise estimation of $e(\bm{\delta})$ is challenging, it is fortunate that only an approximation, denoted as $\tilde{e}(\bm{\delta})$, is sufficient to move the adversarial embedding back within the manifold. As can be observed in Figure 2(e), although the adversarial image after the blur operation does not fall exactly onto the original image, this favorable direction is sufficient to serve as a proxy target to guide the adversarial image correctly onto the natural manifold, as shown in Figure 2(c)(d). We will elaborate on this novel purification pipeline ZeroPur in the following two Sections.
Figure 2: The impact of adversarial samples and their improvement through different operations, visualized using t-SNE. (a) AutoAttack [4] causes adversarial samples to deviate from the manifold; (b) Gaussian Blur, (c) Guided Shift and (d) ZeroPur (Ours) reverse this deviation to varying degrees. (e) Feature trajectories under adversarial attack and the blur operation, indicating that simple blurring operator can provide a promising direction toward the natural image manifold.
@@ 3.2 Guided Shift
where $\bm{\delta}_{\mathrm{pfy}}$ and $\epsilon_{\mathrm{pfy}}$ are defined to correspond to $\bm{\delta}$ and $\epsilon$ in Eq.(2) to offset the perturbation. However, even if $\epsilon_{\mathrm{adv}}$ can be considered as a hyperparameter, the ground-truth label $y$ is not accessible. External model-based adversarial purification [56, 33, 25, 43] typically train a purification model to minimize the global $\ell_{\mathrm{pfy}}$. Auxiliary function-based adversarial method [31, 42, 17] tends to design a suitable $\ell_{\mathrm{pfy}}$ to complete purification without relying on external generative models. They need to retrain the classifier or parameterized auxiliary function to effectively cooperate in removing adversarial perturbations, which means these methods all introduce parameters $\Theta$ in $\ell_{\mathrm{pfy}}$:
@@
Based on our assumption in Section 3.1 that adversarial images are outliers of the natural image manifold and the purification process can be considered as returning them to this manifold, we aim to design $\ell_{\mathrm{pfy}}$ without $\Theta$ by allowing them to shift towards the manifold adaptively:
@@
We are starting to investigate whether a simple image transformation, such as color jitter, grayscale, Gaussian blur, solarization, and equalization, can shift adversarial images towards the natural image manifold. Fortunately, as shown in Figure 2(e), the blurring operation nearly provides a genuine direction toward the natural samples for the adversarial examples. This phenomenon suggests that a simple blurring operator can bring adversarial examples closer to the manifold.
@@
where $\mathrm{blur}(\cdot)$ is the blurring operator and $\bm{x}_{\mathrm{g}}^{0}:=\bm{x}_{\mathrm{adv}}$. The workflow of GS is shown in Algorithm 1, where we use $\mathrm{sgn}(\cdot)$ to regulate the step size.
To demonstrate the process of GS, we employ two blurring operators: the Median filter and Gaussian blur to compute the cosine similarity between single-step result $\bm{x}_{\mathrm{g}}^{t}$ of GS and the natural image $\bm{x}$ ($\star$-$GS$-$N$), as well as between the blurred counterpart $\bm{\tilde{x}}_{\mathrm{g}}^{t}$ of $\bm{x}_{\mathrm{g}}^{t}$ and $\bm{x}$ ($\star$-$GS^{\prime}$-$N$), as shown in Figure 3. Interestingly, as GS progresses, both cosine similarities consistently increase, overcoming the inherent limitations of a single blur. Its effectiveness can be qualitatively described by Figure 2(c), where adversarial examples, guided by their blurred counterparts, gradually approach the natural manifold.
@@ 3.3 Adaptive Projection
We can also observe an interesting phenomenon in Figure 3. Regardless of the type of blurring operators, dash lines ($\star$-$GS^{\prime}$-$N$) and solid lines ($\star$-$GS$-$N$) tend to converge. Therefore, we can have the following assumption:
@@
Figure 4: Intuitive explanation of AP in the context of adversarial or natural $\bm{x}_{\mathrm{init}}$. PR: the perceptual regularization (Eq.(20)).
@@
$\mathcal{L}_{l}(\bm{x}_{\mathrm{init}},\bm{x}_{\mathrm{g}},\bm{x}_{\mathrm{p}})=\Delta\bm{u}^{l}_{\mathrm{p}}\cdot\Delta\bm{u}^{l}_{\mathrm{g}},$ (16)
@@
where $f_{l}$ denotes feature maps at layer $l$ of the classifier, $\bm{x}_{\mathrm{p}}$ is initialized by $\bm{x}_{\mathrm{init}}$. Maximizing Eq.(16) is equivalent to maximizing the projection of $\bm{u}^{l}_{\mathrm{p}}$ onto $\bm{u}^{l}_{\mathrm{g}}$ since $\|\bm{u}^{l}_{\mathrm{g}}\|$ is a constant. The increase in projection allows $\bm{x}_{\mathrm{p}}$ to move freely along the direction of GS, enabling it to independently approach the natural image manifold.
However, We typically do not know whether an image is adversarial in real-world applications (i.e., $\bm{x}_{\mathrm{init}}=\bm{x}$ when $\bm{x}$ is natural), repeatedly guiding images by blurred counterparts may cause natural images to deviate from the manifold. As shown in Figure 4, this deviation will be enhanced by AP. Therefore, we introduce a perceptual regularization term in Eq.(15). Let $F_{1}$ is the dynamic of Eq.(16):
$F_{1}=\lambda_{1}\cdot\nabla_{\bm{x}_{\mathrm{p}}}\frac{1}{\|S\|}\sum_{l\in S}\mathcal{L}_{l}(\bm{x}_{\mathrm{init}},\bm{x}_{\mathrm{g}},\bm{x}_{\mathrm{p}})\ ,$ (18)
@@ 4 Experiments
Datasets and Metrics. Three benchmarks CIFAR-10 [21], CIFAR-100 [21], and ImageNet-1K [8] are considered to evaluate our method. We compare our method with the state-of-the-art adversarial training methods reported in standard benchmark RobustBench [5], as well as with other representative adversarial training and purification methods [42, 31, 26, 6, 22, 46, 15, 56, 48, 33, 25, 27, 43]. In all experiments, we consider two metrics to evaluate the performance of all methods: standard accuracy (SA) and robust accuracy (RA). The standard accuracy measures the performance of the defense method on natural images, while the robust accuracy measures the performance on adversarial images.
Adversarial attacks and victim classifiers. We evaluate our methods with three attacks: AutoAttack [4], DI2-FGSM [54], and BPDA [2]. AutoAttack is a powerful adaptive attack commonly used in most defense studies. We also use DI2-FGSM to seek whether an attack robust to the blurring operator can affect our method. Additionally, we consider BPDA to attack our purification module in the worst-case scenario. These attacks will attack three victim classifiers, including ResNet-18 [13], ResNet-50 [13], and WideResNet-28-10 [58]. For more information about these attacks, refer to Appendix A.3.
Table 1: Comparison with AT &\&& ABP methods under AutoAttack $\ell_{\infty}$ ($\epsilon=8/255$). $\dagger$ indicates this method uses WideResNet-34-10 as a classifier.
Type Require Training ResNet-18 WideResNet-28-10
Cls Aux Method SA(%) RA(%) Method SA(%) RA(%)
@@
(Kulkarni and Weng, 2024) 85.18 54.53 (Hwang et al., 2023$\dagger$ 82.22 56.63
@@
(Kulkarni and Weng, 2024) 65.37 32.55 (Hwang et al., 2023$\dagger$ 64.73 32.86